OpenStudy (anonymous):

Verify that quadrilateral BCDE is a rhombus with vertices B(5, -1), C(-2, 0), D( -1, 7) and E(6, 6) by showing that all four sides are equal.

jimthompson5910 (jim_thompson5910):

BCDE has the sides: BC, CD, DE, ED you need to show that all 4 sides are equal in length

jimthompson5910 (jim_thompson5910):

to find the length of any segment, you use the distance formula example: to find the length of BC, you find the distance from point B to point C

jimthompson5910 (jim_thompson5910):

the distance formula is \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

OpenStudy (anonymous):

I'm trying and i just can't get it

jimthompson5910 (jim_thompson5910):

Let's make B the point (x1,y1) and C the point (x2,y2)

jimthompson5910 (jim_thompson5910):

B = (x1,y1) = (5, -1) C = (x2,y2) = (-2,0)

jimthompson5910 (jim_thompson5910):

plug the values into the distance formula to get \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] \[\Large d = \sqrt{\left(-2-5\right)^2+\left(0-(-1)\right)^2}\] \[\Large d = \sqrt{\left(-7\right)^2+\left(1\right)^2}\] \[\Large d = \sqrt{49+1}\] I'll let you finish

OpenStudy (anonymous):

thank you!!!!

jimthompson5910 (jim_thompson5910):

sure thing